OpenStudy (anonymous):

Could someone else check my work? I can't find my mistake. Function : x^2 y + x y^2 = 2x two points of two tangent lines (1,1) and (1,-2)

OpenStudy (anonymous):

First I find the derivative of x^2 y + x y^2 = 2x which is \[y' = \frac{ 2-2xy-y^2 }{ x^2+2xy }\]

OpenStudy (anonymous):

Next I find the slopes of both tangent lines to form an equation for both lines.

OpenStudy (anonymous):

I plug in (1,1) into y' and get the slope of -1/3. I plug in (1,-2) into y' and get the slope of -2/3

OpenStudy (anonymous):

Now I form an equation with the points and slopes I found.

OpenStudy (anonymous):

y -1 = -1/3 (x-1) y+2 = -2/3 (x-1) y-1 = -1/3x + 1/3 y+2 = -2/3x + 2/3 y = -1/3x + 1/3 + 1 y = -2/3 x + 2/3 - 2 y = -1/3x + 4/3 y = -2/3x - 4/3

OpenStudy (anonymous):

\[y = \frac{ -x + 4 }{ 3 } , y= \frac{- 2x-4 }{3 }\]

OpenStudy (anonymous):

Then I proceed to find x and y where they intersect each other.

OpenStudy (anonymous):

(-x+4)/3 = (-2x-4)/3 3(-x+4) = 3(-2x-4) -3x + 12 = -6x - 12 3 x = -24 x = -8

OpenStudy (anonymous):

y = (-(-8)+4)/3 y = 12/3 y = 4